OpenStudy (anonymous):

The height of an equilateral triangular prism increases by four units. The new volume is more than the original by how much? Answer four more than the the area of the base four more than the length of the height four times the length of the height four times the area of the base

OpenStudy (anonymous):

V1=Sb.h V2=Sb.(h1+4)=Sb.h1+4.Sb-> ->V2=V1+4.Sb

OpenStudy (anonymous):

sb?

OpenStudy (anonymous):

 Volume of prism=1/2 * bhl Volume of new prism = 1/2 *b(h+4)l=1/2 * bhl + 2bl. it's an additional of 2bl.